Let X and Y be two discrete random independent random variables.

p(x) = 1/3 for x =-2,-1,0

p(y) = 1/2 for y =1,6

K = X + Y
Answer #1

MGF of K is given by:

M_{K}(t)=M_{X+Y}(t)=E[e^{t(X+Y)}]=E[e^{tX+tY}]=E[e^{tX}]E[e^{tY}]=M_{X}(t).M_{Y}(t)

Now, M_{X}(t)=sum e^{tx}p(x)=e^{-2t}p(-2)+e^{-t}p(-1)+e^{0t}p(0)

3 2t

M_{Y}(t)=sum e^{ty}p(y)=e^{t}p(1)+e^{6t}p(6)=rac{1}{2}[e^{t}+e^{6t}]

So, M_{K}(t)=rac{1}{3}[e^{-2t}+e^{-t}+1].rac{1}{2}[e^{t}+e^{6t}]=rac{1}{6}[e^{-2t}+e^{-t}+1].[e^{t}+e^{6t}]